OpenStudy (anonymous):

wiggle wiggle What is the equation of the following line in standard form? 11x+y= -7 7x+y= 11 7x+y= -11

OpenStudy (anonymous):

wiggle, YEAH

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

But seriously, we should stop wasting our time.

OpenStudy (anonymous):

ok an pluse i had go get yall to look 11x+y= -7 7x+y= 11 7x+y= -11

myininaya (myininaya):

Hey please don't start your questions off with off-topic-ness. It causes spam. Thanks.

myininaya (myininaya):

I changed your question for you. Thanks.

OpenStudy (anonymous):

ok

OpenStudy (experimentx):

find those two points ... use double point formula http://en.wikipedia.org/wiki/Linear_equation#Two-point_form